Question four:

i)When a new partner was admitted into partnership, he had to pay $5,000 as goodwill.There was a squabble between the senior partner and the bookkeeper. Thebookkeeper’s argument was that the money needs to be credited to the account of newpartner because after all it is his money”. However senior partner strongly disagreed tothis treatment.


Required:

Give your advice as to the proper treatment of this $5,000. Explain your reasons fully.


ii)Teko and Thuto are partners in a firm sharing profits in the ratio of 3:2. They admit Keiso as a new partner. Teko sacrificed 1/5 of his share and Thuto 1/4 of his share in favour of Keiso. Calculate new profit sharing ratio of Teko, Thuto and Keiso.


iii) Assume Teko and Thuto’s share capitals are $40,000 and &30,000 respectively. Keiso will invest $18,000 for one-third interest in capital. Goodwill is recorded.Show the journal entries.


1
Expert's answer
2022-02-17T04:38:08-0500

i) Goodwill should be recorded and allocated to the partners as per the profit/loss sharing ratio

ii)

Teko= "\frac{3}{5}"


Thuo= "\frac{2}{5}"


Keiso= "(\frac{1}{5}\times\frac {3}{5})\times (\frac{1}{4}\times \frac{2}{5})= \frac{1}{50}"


New Ratios

"Teko= \\frac{3}{5}- \\frac{1}{50}= \\frac{29}{50}\n\n\u200b"

"Thuo= \\frac{2}{5}- \\frac{1}{50}= \\frac{19}{50}\n\u200b"

New ratio= Teko: Thuo: Keiso= "\frac{29}{50}: \frac{19}{50}: \frac{1}{50}= 29:19:1"

iii)

Value ofGood will = "\frac{(40,000+30,000+18000)}{3}"

= 29,333

Less Keiso contribution= 29,333- 18000= 11,333

Therefore goodwill recognized by new partners= "\frac{11,333}{\frac{2}{3}}= 17,000"

Journal entries